# Lifelong Learning with Behavior Consolidation for Vehicle Routing

### v2

@@ 1 Introduction
Figure 1: Catastrophic forgetting while fine-tuning a solver on sequential, new tasks with different distributions and scales.
@@
Neural solvers have gained significant attention for their remarkable performance on vehicle routing problems (VRPs) (Vinyals et al., 2015; Bello et al., 2017; Bengio et al., 2021; Bogyrbayeva et al., 2024). By training on one or multiple specific problem distributions and scales, i.e., tasks, simultaneously, neural solvers can learn to construct high-quality solutions or improve existing solutions efficiently for problems drawn from the same distribution and of the same scale as the training tasks (Kool et al., 2019; Kwon et al., 2020; Lu et al., 2020; Tang & Yao, 2024). However, in real-world scenarios, unpredictable problems with new distributions or scales often arise over time. For example, a daily delivery service provider receives different numbers of orders with new patterns of delivery locations and demands due to unpredictable consumer behavior, business expansion, and new shopping trends. In such scenarios, it is challenging to develop a universal solver that can perform effectively across all possible problem distributions and scales after one-off training (Liu et al., 2023). Some recent studies (Jiang et al., 2022; Bi et al., 2022; Gao et al., 2024; Wang et al., 2024b; Xiao et al., 2025) address unpredictable new problems by enhancing the generalization ability of trained solvers. However, the generalization would have a boundary, and on new problems that differ substantially from the training ones, these methods still require additional learning to enhance problem-solving performance (Zhou et al., 2023). A common approach is to fine-tune the trained solver on a new task (Bengio et al., 2021; Zhou et al., 2023); nonetheless, it leads to catastrophic forgetting of knowledge acquired from previous tasks (Khetarpal et al., 2022). Solver’s performance on problems following the previous tasks will decrease significantly (c.f., Figure 1), as the model parameters will be overwritten during training on a new task (Wang et al., 2024c). These issues highlight the need for neural solvers capable of lifelong learning, enabling them to continuously acquire knowledge from unforeseen tasks while retaining generic knowledge and their performance on previous tasks.
@@
To our best knowledge, only Li et al. (2024) and Feng et al. (2025) study lifelong learning for neural VRP solvers. However, the scenarios studied in both works are significantly different from ours: they neglect the possibility of distributional shifts across tasks and primarily focus on a fixed task order (e.g., the order that problem scale constantly increases). Though effective in their focused scenarios, their proposed methods rely on actively generating and training on new problem instances from previous tasks, making them inapplicable when the generation of new instances is uncontrollable, as is the case studied in this work. In contrast, this work focuses on a realistic and general setting, jointly addressing both scale and distribution changes between tasks without requiring any prior knowledge or assumptions about task order. We propose a novel method for lifelong learning of neural VRP solvers, called Lifelong Learning Router with Behavior Consolidation (LLR-BC). LLR-BC focuses on addressing the catastrophic forgetting issue that arises during training across new tasks arising sequentially, which is commonly overlooked in prior neural VRP solver research. It buffers and revisits experiences from previous tasks to retain acquired knowledge. To efficiently utilize the limited experience memory, LLR-BC weights experiences based on decision confidence, emphasizing more crucial ones. Given that small changes in action probability distribution, especially under low confidence, can alter decisions and drastically affect the constructed route, LLR-BC minimizes reverse Kullback–Leibler divergence to consolidate behavior in a decision-seeking way. Extensive experiments on capacitated vehicle routing problems (CVRPs) and traveling salesman problems (TSPs) under various task sequences and base neural solvers show that LLR-BC effectively mitigates forgetting, maintains plasticity, and improves zero-shot generalization by accumulating transferable knowledge over time.
@@
Our major contributions include: i) we introduce LLR-BC, a general lifelong learning framework enabling neural VRP solvers to learn from different tasks sequentially and address catastrophic forgetting effectively; ii) we propose Confidence-aware Experience Weighting (CaEW) to prioritize crucial experiences, as a module of LLR-BC for enhancing the effectiveness of experiences utilizing; iii) we propose Decision-seeking Behavior Consolidation (DsBC), which works in LLR-BC to preserve past behaviors by minimizing the discrepancy between buffered and current solver behaviors on stored states, with an emphasis on replicating past decisions; iv) the effectiveness of LLR-BC in is validated through extensive experiments, supported by detailed analysis and discussion.
@@ 2 Related Work
Machine learning, particularly deep reinforcement learning (DRL), enables neural solvers to learn VRP-solving strategies directly from problem-solving experience, eliminating the need for hand-crafted heuristics (Bengio et al., 2021). Existing methods are categorized as neural construction or neural improvement approaches (Bi et al., 2022; Kong et al., 2024). Neural construction methods build solutions from scratch by sequentially selecting the next node in an autoregressive fashion. Notable examples such as POMO (Kwon et al., 2020) and DualOpt (Zhou et al., 2025) generate high-quality solutions within seconds, rivaling strong heuristics like LKH3 (Helsgaun, 2017). In contrast, neural improvement methods learn to enhance existing solutions, either by configuring (Wu et al., 2022; Ma et al., 2023) or selecting (Lu et al., 2020; Pei et al., 2025b; Chen et al., 2025; Guo et al., 2025) among predefined improvement heuristics. Although improvement methods have the potential to further improve solution quality by increasing the number of iterations, this also leads to higher computational cost. This paper focuses on lifelong learning for neural construction methods, due to their favorable balance between effectiveness and efficiency.
@@ 2.2 Cross-distribution Cross-scale Generalization
Early studies of neural VRP solvers typically focus on solving a single task, i.e., problems with fixed scale and one given distribution (e.g., uniform) (Kool et al., 2019; Kwon et al., 2020), resulting in poor generalization across distributions and scales (Bi et al., 2022; Xiao et al., 2025), which is often encountered in real-world applications. To address this issue, recent studies focus on improving cross-distribution or cross-scale performance. DROP (Jiang et al., 2022) improves cross-distribution performance by training across multiple distributions and optimizing for worst-case task performance. AMDKD (Bi et al., 2022) distills knowledge from multiple task-specific teacher models into a student model, achieving improvement in zero-shot generalization. ELG (Gao et al., 2024) separates local and global policies and, despite training only on uniform problems of scale 100, generalizes well across other tasks. Xiao et al. (2025) propose two network modules to handle different scales and distributions, respectively, improving zero- and few-shot generalization. INViT (Fang et al., 2024) involves multiple encoders and decoders with multiple views to improve zero-shot generalization. Omni (Zhou et al., 2023) applies meta-learning to learn a solver that can quickly adapt to a new task.
However, these works mainly focus on one-off training on a pre-defined task set. New tasks are addressed with zero-shot generalization or fine-tuning the trained solver on each of them independently. The potential of sequential knowledge accumulation from multiple new tasks is overlooked, together with the impact of catastrophic forgetting. As a result, their performance in learning to solve sequentially arising tasks is still limited. In practice, tasks arrive continuously, and past tasks can inform future learning. This motivates the need for lifelong neural VRP solvers that can incrementally learn from a sequence of diverse tasks.
@@ 2.3 Lifelong Learning
Several strategies have been proposed to mitigate catastrophic forgetting. Training task-specific model components is an example (Wu et al., 2021; Ebrahimi et al., 2020), but this increases storage costs and poses challenges in selecting the correct model when the task of the test data is unknown. Regularization on parameter updating to preserve prior knowledge is also widely used (Kirkpatrick et al., 2017; Zenke et al., 2017). However, this can hinder learning on new tasks. A further widely adopted method is experience replay, where past experiences on previously learned tasks are stored and revisited during training on a new task (Isele & Cosgun, 2018; Buzzega et al., 2020). Nonetheless, identifying and retaining the most crucial experiences to improve memory efficiency and minimize interference with learning new tasks often requires problem-specific designs.
@@ 3 Lifelong Learning Router with Behavior Consolidation
Figure 2: LLR-BC in the lifelong learning scenario where new tasks with different distributions and scales sequentially arise over time.
We consider the scenarios where only new problem instances of the current task are generated for training (i.e., ), while future tasks and their order are unknown, reflecting practical cases. We assume that all the tasks are equally important, and each has a fixed and known training budget. The main aim is to train a solver capable of generating high-quality solutions across all the learned tasks arising over time. So, a trained solver is evaluated on all the tasks learned so far.
We propose LLR-BC, inspired by the experience replay paradigm (Isele & Cosgun, 2018; Buzzega et al., 2020). It consolidates knowledge learned from previous tasks by revisiting the old experiences obtained from these tasks. To increase the effectiveness of utilizing limited old experiences, we introduce two ingredients in LLR-BC: CaEW and DsBC. Figure 2 illustrates LLR-BC’s full workflow. LLR-BC maintains a fixed-size experience memory and trains the solver learned from the previous tasks on each newly arrived task. Following standard practice in neural VRP solvers, LLR-BC trains on each task in epochs. In each epoch, LLR-BC iteratively samples and solves a batch of problem instances $\mathcal{I}$ from the current task to obtain a set of experience trajectories $\{\tau\}$. A DRL algorithm that maximizes the reward gain based on $\{\tau\}$ is applied to update the solver model. Simultaneously, a subset of experiences $\mathcal{E}$ is sampled from the memory for consolidation of previously learned tasks. CaEW assigns higher weights to more crucial sampled experiences, then DsBC guides model updates by minimizing weighted behavioral divergence between the current model and the behavior buffer of the sampled experiences. This joint process mitigates catastrophic forgetting while preserving adaptability. Notably, LLR-BC is a general and readily integrable framework, independent of specific model architectures or RL algorithms, and applicable to many existing neural solvers, such as POMO (Kwon et al., 2020) and INViT (Fang et al., 2024). Core ingredients of LLR-BC are described below. More details about LLR-BC can be found in Appendix A.
@@ 3.1 Experience Replay for Routing
We follow the commonly studied Markov decision process (MDP) formulation of constructive neural VRP solver (Kwon et al., 2020; Zhou et al., 2023; Bogyrbayeva et al., 2024), where each state corresponds to a partial solution composed of visited nodes, and an action is to select one node to visit next. Each experience $e=\langle s,\mathcal{P}\rangle$ consists of a state $s$ and the solver’s behavior $\mathcal{P}$ on the state $s$. A state $s$ represents the current partial solution for solving a problem. The behavior $\mathcal{P}$ is the probability distribution of selecting each node in a problem as the next visit. The detailed solver design, including the handling of infeasible actions and the encoding of problems and solutions, follows any underlying base neural solver used. We use the probability distribution over all nodes rather than the single selected node to represent the solver’s behavior, as it captures richer information about the learned routing strategy (Rusu et al., 2016; Buzzega et al., 2020).
To consolidate high-quality behaviors, LLR-BC buffers experiences only during the final epoch of each task. It employs reservoir sampling (Vitter, 1985) to maintain a balanced and fixed-size buffer $\mathcal{B}$. New experiences are added directly if the memory is not full; otherwise, each incoming experience replaces a randomly selected experience entry with probability $\frac{|\mathcal{B}|}{N}$, where $|\mathcal{B}|$ is the buffer size and $N$ denotes the sum of the number of experiences obtained from the final epoch of training on each previously learned task, i.e., the total number of experiences considered for buffering across all encountered tasks so far. Reservoir sampling ensures that all experiences are equally likely to be buffered while the total number of obtained experiences is increasing. At each model update step, a random set of experiences $\mathcal{E}$ is sampled from buffer $\mathcal{B}$ for behavior consolidation. A detailed discussion of the additional memory usage from the buffer can be found in Appendix A.4.
@@ 3.2 Confidence-aware Experience Weighting (CaEW)
Decisions made with low confidence could be more susceptible to change during model updates (Farahmand, 2011; Ahmed et al., 2019). Therefore, LLR-BC assigns higher consolidation weights to such experiences, encouraging the solver to preserve behaviors in crucial decision points. Confidence is estimated by the variance of the action probability distribution, where lower variance indicates lower confidence and leads to greater emphasis during consolidation. The normalized confidence-based weight of an experience $e=\langle s,\mathcal{P}\rangle$ is computed as $w(e)=1-\frac{\text{var}(\mathcal{P})}{\text{var}_{\text{max}}(|\mathcal{P}|)}$, where $\text{var}(\mathcal{P})$ is the variance of $\mathcal{P}$ and $|\mathcal{P}|$ is the size of $\mathcal{P}$, i.e., the number of actions. $\text{var}_{\text{max}}(|\mathcal{P}|)=\frac{|\mathcal{P}|-1}{|\mathcal{P}|^{2}}$ denotes the maximum possible variance for a distribution with the same number of candidates as $\mathcal{P}$. Then, given a set of sampled experiences $\mathcal{E}$, we rescale weights such that the sum of weight equals to 1, i.e., $\bar{w}_{\mathcal{E}}(e)=\frac{w(e)}{\sum_{e^{\prime}\in\mathcal{E}}w(e^{\prime})}$, $\forall e\in\mathcal{E}$.

### v3

@@ 1 Introduction
Figure 1: Conceptual demonstration of catastrophic forgetting while fine-tuning on sequential, new tasks with different distributions and scales.
@@
Neural solvers have gained significant attention for their remarkable performance on vehicle routing problems (VRPs) (Vinyals et al., 2015; Bello et al., 2017; Bengio et al., 2021; Bogyrbayeva et al., 2024). By training on one or multiple specific problem distributions and scales, i.e., tasks, simultaneously, neural solvers can learn to construct high-quality solutions or improve existing solutions efficiently for problems drawn from the same distribution and of the same scale as the training tasks (Kool et al., 2019; Kwon et al., 2020; Lu et al., 2020; Tang and Yao, 2024). However, in real-world scenarios, unpredictable problems with new distributions or scales often arise over time. For example, a daily delivery service provider receives different numbers of orders with new patterns of delivery locations and demands due to unpredictable consumer behavior, business expansion, and new shopping trends. In such scenarios, it is challenging to develop a universal solver that can perform effectively across all possible problem distributions and scales after one-off training (Liu et al., 2023). Some recent studies (Jiang et al., 2022; Bi et al., 2022; Gao et al., 2024; Wang et al., 2024b; Xiao et al., 2025) address unpredictable new problems by enhancing the generalization ability of trained solvers. However, the generalization would have a boundary, and on new problems that differ substantially from the training ones, these methods still require additional learning to enhance problem-solving performance (Zhou et al., 2023). A common approach is to fine-tune the trained solver on a new task (Bengio et al., 2021; Zhou et al., 2023); nonetheless, it leads to catastrophic forgetting of knowledge acquired from previous tasks (Khetarpal et al., 2022). Solver’s performance on problem instances of the previous tasks will decrease significantly (cf. Figure 1), as the model parameters will be overwritten during training on a new task (Wang et al., 2024c). These issues highlight the need for neural solvers capable of lifelong learning, enabling them to continuously acquire knowledge from unforeseen tasks while retaining generic knowledge and their performance on previous tasks.
Existing studies on lifelong learning for neural VRP solvers (Li et al., 2025; Feng et al., 2025) are largely confined to highly specific scenario settings: tasks solely differing in scale or distance metric, the task order is fixed and known, and the generation of new problem instances is controllable. Furthermore, they are mainly dedicated to the solver’s performance at the end of the lifelong learning process after being trained on all training tasks. As lifelong learning is an ongoing process (possibly without any end), the solver’s performance at each time point of the lifelong learning process is also important. Their proposed methods rely on actively generating and training on new problem instances from previous tasks, making them inapplicable when the generation of new instances is uncontrollable, as is the case studied in this work.
@@
In contrast, this work targets more realistic and general lifelong learning scenarios, simultaneously accommodating changes in both scale and distribution across tasks, without assuming any prior knowledge of the task order or any control over problem instance generation. We propose a novel method for lifelong learning of neural VRP solvers, called Lifelong Learning Router with Behavior Consolidation (LLR-BC). LLR-BC focuses on addressing the catastrophic forgetting issue that arises during training across new tasks arising sequentially, which is commonly overlooked in prior neural VRP solver research. It buffers and revisits experiences from previous tasks to retain acquired knowledge. To efficiently utilize the limited experience memory, LLR-BC weights experiences based on decision confidence, emphasizing more crucial ones. Given that small changes in action probability distribution, especially under low confidence, can alter decisions and drastically affect the constructed route, LLR-BC minimizes reverse Kullback–Leibler divergence to consolidate behavior in a decision-seeking way. Extensive experiments on capacitated vehicle routing problems (CVRPs) and traveling salesman problems (TSPs), equipped with a comprehensive metrics set to evaluate the solver during the ongoing lifelong learning process, under various task sequences and base neural solvers show that LLR-BC effectively mitigates forgetting, maintains plasticity, and improves zero-shot generalization by accumulating transferable knowledge over time.
@@
Our major contributions include: i) we introduce LLR-BC, a general lifelong learning framework enabling neural VRP solvers to learn from different tasks sequentially and address catastrophic forgetting effectively; ii) we propose Confidence-aware Experience Weighting (CaEW) to prioritize crucial experiences, as a module of LLR-BC for enhancing the effectiveness of experiences utilization; iii) we propose Decision-seeking Behavior Consolidation (DsBC), which works in LLR-BC to preserve past behaviors by minimizing the discrepancy between buffered and current solver behaviors on stored states, with an emphasis on replicating past decisions; iv) the effectiveness of LLR-BC in is validated through extensive experiments, supported by detailed analysis and discussion.
@@ 2.1 Neural VRP Solvers
Machine learning, particularly deep reinforcement learning (DRL), enables neural solvers to learn VRP-solving strategies directly from problem-solving experience, eliminating the need for hand-crafted heuristics (Bengio et al., 2021). Existing methods are categorized as neural construction or neural improvement approaches (Bi et al., 2022; Kong et al., 2024). Neural construction methods build solutions from scratch by sequentially selecting the next node in an autoregressive fashion. Notable examples such as POMO (Kwon et al., 2020) and DualOpt (Zhou et al., 2025) generate high-quality solutions within seconds, rivaling strong heuristics like LKH3 (Helsgaun, 2017). In contrast, neural improvement methods learn to enhance existing solutions, either by configuring (Wu et al., 2022; Ma et al., 2023) or selecting (Lu et al., 2020; Pei et al., 2025a; Chen et al., 2025; Guo et al., 2025) among predefined improvement heuristics. Although improvement methods have the potential to further improve solution quality by increasing the number of iterations, this also leads to higher computational cost. This paper focuses on lifelong learning for neural construction methods, due to their favorable balance between effectiveness and efficiency.
@@ 2.2 Cross-distribution Cross-scale Generalization
Early studies of neural VRP solvers typically focus on solving a single task, i.e., problems with fixed scale and one given distribution (e.g., uniform) (Kool et al., 2019; Kwon et al., 2020), resulting in poor generalization across distributions and scales (Bi et al., 2022; Xiao et al., 2025), which is often encountered in real-world applications. To address this issue, recent studies focus on improving cross-distribution or cross-scale performance. DROP (Jiang et al., 2022) improves cross-distribution performance by training across multiple distributions and optimizing for worst-case task performance. AMDKD (Bi et al., 2022) distills knowledge from multiple task-specific teacher models into a student model, achieving improvement in zero-shot generalization. ELG (Gao et al., 2024) separates local and global policies and, despite training only on uniform problems of scale 100, generalizes well across other tasks. Xiao et al. (2025) proposes two network modules to handle different scales and distributions, respectively, improving generalization. INViT (Fang et al., 2024) involves multiple encoders and decoders with multiple views to improve generalization. Omni (Zhou et al., 2023) applies meta-learning to learn a solver that can quickly adapt to a new task.
However, these works mainly focus on one-off training on a predefined task set. New tasks are addressed with zero-shot generalization or fine-tuning the trained solver on each of them independently. The potential of sequential knowledge accumulation from multiple new tasks is overlooked, together with the impact of catastrophic forgetting. As a result, their performance in learning to solve sequentially arising tasks is still limited. In practice, tasks arrive continuously, and past tasks can inform future learning. This motivates the need for lifelong neural VRP solvers that can incrementally learn from a sequence of diverse tasks.
@@ 2.3 Lifelong Learning
Several strategies have been proposed to mitigate catastrophic forgetting. Training task-specific model components is an example (Wu et al., 2021; Ebrahimi et al., 2020), but this increases storage costs and poses challenges in selecting the correct model when the task of the test data is unknown. Regularization of parameter updating to preserve prior knowledge is also widely used (Kirkpatrick et al., 2017; Zenke et al., 2017). However, this can hinder learning on new tasks. A further widely adopted method is experience replay, where past experiences on previously learned tasks are stored and revisited during training on a new task (Isele and Cosgun, 2018; Buzzega et al., 2020). Nonetheless, identifying and retaining the most crucial experiences to improve memory efficiency and minimize interference with learning new tasks often requires problem-specific designs.
To the best of our knowledge, only two existing works have studied lifelong learning for neural VRP solver (Li et al., 2025; Feng et al., 2025). However, their studies focused on rather restricted scenarios: (i) tasks differ only in scale (Li et al., 2025; Feng et al., 2025) or distance metric (Feng et al., 2025), (ii) task order is fixed and known under strong assumptions (e.g., task scale gradually increases), and (iii) problem instance generation is controllable so that new instances of previously learned tasks can always be actively generated, making these methods inapplicable to scenarios with uncontrollable instance streams. Studies of more generic and practical scenarios are desired.
@@ 3 Lifelong Learning Router with Behavior Consolidation
Figure 2: LLR-BC in the lifelong learning scenario where new tasks with different distributions and scales sequentially arise over time. $T_{t}$: the task at time $t$. $\pi_{\theta}$: solver with model parameters $\theta$. $\mathcal{B}$: experience buffer. $\mathcal{J}$: generated problem instances. $\{\tau\}$: problem solving trajectories. $\mathcal{E}$: experiences sampled from the buffer. $w_{\mathcal{E}}$: weights of experiences. $a$: action from action space $\mathcal{A}$, i.e., node to visit. $s$: state. var: variance. $\mathcal{P}$: probability distribution over actions.
We consider the scenarios where problem instance generation is uncontrollable and only new instances of the current task are generated for training, while future tasks and their order are unknown, reflecting practical cases. We assume that all the tasks are equally important, and each has a fixed and known training budget. The main aim is to train a solver capable of generating high-quality solutions across all the learned tasks arising over time. Therefore, after training on each new task, the solver is tested on all tasks learned so far.
We propose LLR-BC, inspired by the experience replay paradigm (Isele and Cosgun, 2018; Buzzega et al., 2020). It consolidates knowledge learned from previous tasks by revisiting the old experiences obtained from these tasks. To increase the effectiveness of utilizing limited old experiences, we introduce two ingredients in LLR-BC: CaEW and DsBC. Figure 2 illustrates LLR-BC’s full workflow. LLR-BC maintains a fixed-size experience memory and trains the solver learned from the previous tasks on each newly arrived task. Following standard practice in neural VRP solvers, LLR-BC trains on each task in epochs. In each epoch, LLR-BC iteratively samples and solves a batch of problem instances $\mathcal{I}$ from the current task to obtain a set of experience trajectories $\{\tau\}$. A DRL algorithm that maximizes the reward gain based on $\{\tau\}$ is applied to update the solver model. Simultaneously, a subset of experiences $\mathcal{E}$ is sampled from the memory for consolidation of previously learned tasks. CaEW assigns higher weights to more crucial sampled experiences, then DsBC guides model updates by minimizing weighted behavioral divergence between the current model and the behavior buffer of the sampled experiences. This joint process mitigates catastrophic forgetting while preserving adaptability. Notably, LLR-BC is a general and readily integrable framework, independent of specific model architectures or RL algorithms, and applicable to many existing neural solvers, such as POMO (Kwon et al., 2020) and INViT (Fang et al., 2024). Core ingredients of LLR-BC are described below. More details about LLR-BC can be found in Appendix A.
@@ 3.1 Experience Replay for Routing
LLR-BC employs reservoir sampling (Vitter, 1985) to maintain a fixed-size buffer $\mathcal{B}$. New experiences are added directly if the memory is not full. Otherwise, each incoming experience replaces a randomly selected buffered experience with probability $\frac{|\mathcal{B}|}{N}$, where $|\mathcal{B}|$ is the buffer size and $N$ denotes the total number of experiences tried to add so far. Reservoir sampling ensures that all experiences are equally likely to be buffered while the total number of obtained experiences is increasing.
While existing methods (Li et al., 2025; Feng et al., 2025) take one whole problem instance of previous tasks as an experience, we consider experiences at finer granularity. We follow the commonly studied Markov decision process (MDP) formulation of constructive neural VRP solver (Kwon et al., 2020; Zhou et al., 2023; Bogyrbayeva et al., 2024), where each state corresponds to a partial solution composed of visited nodes, and an action is to select one node to visit next. In LLR-BC, each experience $e=\langle s,\mathcal{P}\rangle$ consists of a state $s$ and the solver’s behavior $\mathcal{P}$ on the state $s$. A state $s$ represents the current partial solution for solving an instance. The behavior $\mathcal{P}$ is the probability distribution of selecting each node as the next visit. The detailed solver design, including the handling of infeasible actions and the encoding of problems and solutions, follows any underlying base neural solver used. We use the probability distribution over all nodes rather than the single selected node to represent the solver’s behavior, as it captures richer information about the learned routing strategy (Rusu et al., 2016; Buzzega et al., 2020).
To consolidate high-quality behaviors, LLR-BC buffers experiences only during the final epoch of each task, as in the final epoch the solver is well trained on the task and can provide high-quality behaviors. At each model update step, a random set of experiences $\mathcal{E}$ is sampled from buffer $\mathcal{B}$ for behavior consolidation. A detailed discussion of the additional memory usage from the buffer can be found in Appendix A.5.
@@ 3.2 Confidence-aware Experience Weighting (CaEW)
Decisions made with low confidence could be more susceptible to change during model updates (Farahmand, 2011; Ahmed et al., 2019). Therefore, LLR-BC assigns higher consolidation weights to such experiences, encouraging the solver to preserve behaviors in crucial decision points. Confidence is measured by the variance of the action probability distribution, where lower variance indicates lower confidence and leads to greater emphasis during consolidation (Spielberg and Azaria, 2019; Balasuntharam et al., 2023). Confidence-aware weight of an experience $e=\langle s,\mathcal{P}\rangle$ is normalized by $w(e)=1-\frac{\text{var}(\mathcal{P})}{\text{var}_{\text{max}}(|\mathcal{P}|)}$, where $\text{var}(\mathcal{P})$ is the variance of $\mathcal{P}$ and $|\mathcal{P}|$ is the size of $\mathcal{P}$, i.e., the number of actions. $\text{var}_{\text{max}}(|\mathcal{P}|)=\frac{|\mathcal{P}|-1}{|\mathcal{P}|^{2}}$ denotes the maximum possible variance for a distribution with the same number of candidates as $\mathcal{P}$. Then, given a set of sampled experiences $\mathcal{E}$, we rescale weights such that the sum of weight equals to 1, i.e., $\bar{w}_{\mathcal{E}}(e)=\frac{w(e)}{\sum_{e^{\prime}\in\mathcal{E}}w(e^{\prime})}$, $\forall e\in\mathcal{E}$. Appendix A.4 provides further discussion about the design of CaEW.